A rectangular container measuring 24 cm 48 cm 55 cm is filled with water. What is the mass of this volume of water in kilograms?
Answer #1

Volume of the container, v = (24*48*55) cc = 63360 cc

Density of water , d = 1gm/cc

Mass of the water in the container, m = v*d = (63360*1 ) gm

= 63360 gm = (63360/1000) = 63.36 kg (ans)